OpenStudy (anonymous):

someone help me please <3 <3 <3

OpenStudy (anonymous):

x+y=-2 2x-3y=-9 Which is the solution of the system

OpenStudy (anonymous):

mutiple the first equation with 3 and then add the second equation.

OpenStudy (anonymous):

USE SUBSTITUTION METHOD

OpenStudy (anonymous):

Multiplying the first equation and adding it to the second: 3x + 3y = -6 3x + 3y = -6 2x - 3y = -9 5x = -15 so, x = -3 Substituting that into the first equation: 3(-3) + 3y = -6 -> -9 + 3y = -6 -> 3y = 9 - 6 = 3 y = 1

OpenStudy (anonymous):

The solution is the ordered pair: (-3, 1)

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome! Good luck to you in you studies, @mariisol

OpenStudy (anonymous):

ilove you

OpenStudy (anonymous):

You made my day. :-)

OpenStudy (anonymous):

thx for the recognition! @mariisol